### v1

# FRRI: a novel algorithm for fuzzy-rough rule induction

Abstract
Interpretability is the next frontier in machine learning research. In the search for white box models — as opposed to black box models, like random forests or neural networks — rule induction algorithms are a logical and promising option, since the rules can easily be understood by humans. Fuzzy and rough set theory have been successfully applied to this archetype, almost always separately. As both approaches to rule induction involve granular computing based on the concept of equivalence classes, it is natural to combine them. The QuickRulesJensenCornelis2009 algorithm was a first attempt at using fuzzy rough set theory for rule induction. It is based on QuickReduct, a greedy algorithm for building decision reducts. QuickRules already showed an improvement over other rule induction methods. However, to evaluate the full potential of a fuzzy rough rule induction algorithm, one needs to start from the foundations. In this paper, we introduce a novel rule induction algorithm called Fuzzy Rough Rule Induction (FRRI). We provide background and explain the workings of our algorithm. Furthermore, we perform a computational experiment to evaluate the performance of our algorithm and compare it to other state-of-the-art rule induction approaches. We find that our algorithm is more accurate while creating small rulesets consisting of relatively short rules. We end the paper by outlining some directions for future work.
keywords:
fuzzy rough set theory, rule induction, classification
[…]
1 Introduction
Many fuzzy rule induction algorithms have been established, mostly for deriving a concise set of rules comprehensible by humans for tasks like classification and prediction. These include, for example, fuzzy association rule mining Dong2005 ; 1593646 , first-order fuzzy rule generation PradeRIchardSerr2003 ; DROBICS2003131 , and linguistic semantics-preserving modeling MarnBlzquez2002FromAT ; QIN2008435 . However, the efficacy of most of the existing approaches to fuzzy rule induction is reduced as the data dimensionality increases. Some methods manage to avoid this, for example standard covering algorithms for rule induction (e.g., RIPPER COHEN1995115 and its fuzzy counterpart, FURIA FURIA ) that learn rules in an incremental way, with each rule in turn constructed by adding maximally informative features one by one.
QuickRules JensenCornelis2009 was developed to combat this problem by directly integrating the rule induction process in a greedy feature selection algorithm based on fuzzy rough set theory fuzzyroughsets ; jensen2008computational . Despite promising initial results reported in JensenCornelis2009 , the greedy nature of this algorithm does not allow it to outperform state-of-the-art algorithms w.r.t. balanced accuracy, ruleset size and rule length, as we will also demonstrate in the experimental section of this paper.
In this paper, we propose a novel rule induction algorithm called Fuzzy-Rough Rule Induction (FRRI). FRRI combines the best ingredients of fuzzy rule induction algorithms and rough rule induction algorithms to induce small sets of short rules that accurately summarize the training data and predict classes of new objects.
The remainder of this paper is structured as follows. Section 2 recalls the required theoretical background, including a short introduction on fuzzy set theory and rough set theory, and examines the existing work on fuzzy and rough rule induction. We give the explanation of the novel FRRI algorithm in Section 3. Section 4 describes the experiments that we set up to evaluate the performance of FRRI and to compare it to state-of-the-art fuzzy and rough set rule induction algorithms, in terms of prediction accuracy, number of generated rules and rule length. Finally, Section 5 concludes the paper and outlines our future work.
2 Preliminaries
2.1 Fuzzy sets
In this paper, $U$ will indicate a finite, non-empty set, called the universe. A fuzzy set fuzzysets $A$ in $U$ is a function $A\colon U\to[0,1]$. For $x$ in $U$, $A(x)$ is called the membership degree of $x$ in $A$. The set of all fuzzy sets of $U$ is denoted as $\mathscr{F}(U)$. Let $A,B\in\mathscr{F}(U)$ be two fuzzy sets. $A$ is a fuzzy subset of $B$ if
$(\forall u\in U)(A(u)\leq B(u))$
The union of a fuzzy set $A$ and a fuzzy set $B$ is the fuzzy set
$(\forall u\in U)((A\cup B)(u)=\max(A(u),B(u))$
In fuzzy logic and fuzzy set theory, we use triangular norms and implicators to generalize the classical connectives conjunction ($\wedge$) and implication ($\rightarrow$), respectively. A triangular norm or t-norm is a function $\operatorname{\mathcal{T}}\colon[0,1]^{2}\to[0,1]$ which has $1$ as neutral element, is commutative and associative and is increasing in both arguments. An implicator is a function $\operatorname{\mathcal{I}}\colon[0,1]^{2}\to[0,1]$ which is decreasing in the first argument and increasing in the second one. Additionally, the following boundary conditions should hold:
(0,0)=(0,1)=(1,1)=1 and $\operatorname{\mathcal{I}}(0,0)=\operatorname{\mathcal{I}}(0,1)=\operatorname{\mathcal{I}}(1,1)=1\text{ and }\operatorname{\mathcal{I}}(1,0)=0.$
A (binary) fuzzy relation on a universe $U$ is simply a fuzzy subset of $U\times U$. A fuzzy relation $R$ is reflexive if $R(x,x)=1$ for all elements $x\in U$. It is symmetric if for every pair of elements $x$, $y$ in the universe, $R(x,y)=R(y,x)$. For a given t-norm $\operatorname{\mathcal{T}}$, $R$ is $\operatorname{\mathcal{T}}$-transitive if for every triplet of elements $x,y,z$, $\operatorname{\mathcal{T}}(R(x,y),R(y,z))\leq R(x,z)$. A reflexive and $\operatorname{\mathcal{T}}$-transitive relation is called a $\operatorname{\mathcal{T}}$-preorder relation. A fuzzy tolerance relation is a reflexive and symmetric fuzzy relation. Given an element $x\in U$, its fuzzy tolerance class w.r.t. to a tolerance relation $R$ is the fuzzy set $Rx(y)=R(x,y)$, for any $y\in U$.
2.2 Fuzzy rough sets
In classical rough set theory roughsets , we examine a universe divided into equivalence classes by an equivalence relation. We can then approximate a concept (i.e., a subset of $U$) using two crisp sets: the lower and the upper approximation. The lower approximation contains all elements which are certainly part of the concept. It is defined as the union of the equivalence classes which are subsets of the concept. The upper approximation, which contains all elements which might possibly be a part of the concept, is defined as the union of all equivalence classes which have a non-empty intersection with the concept in question. In fuzzy rough sets fuzzyroughsets , we replace the crisp equivalence relation with a binary fuzzy relation $R$ on $U$ and, in turn, the upper and lower approximations also become fuzzy sets. Let $A$ be a fuzzy set in $U$. Furthermore, let $\operatorname{\mathcal{T}}$ be a t-norm and $\operatorname{\mathcal{I}}$ be an implicator. We define the fuzzy rough set (FRS) for the fuzzy set $A$ w.r.t. the fuzzy relation $R$ as the pair $(\underline{A}_{R}^{\operatorname{\mathcal{I}}},\overline{A}_{R}^{\operatorname{\mathcal{T}}})$, with, for $x$ in $U$:
  • The fuzzy lower approximation $\underline{A}_{R}^{\operatorname{\mathcal{I}}}$ of $A$ is the fuzzy set
    $\underline{A}_{R}^{\operatorname{\mathcal{I}}}(x)=\min_{y\in U}\operatorname{\mathcal{I}}(R(x,y),A(y))$ (1)
  • The fuzzy upper approximation $\overline{A}_{R}^{\operatorname{\mathcal{T}}}$ is the fuzzy set
    $\overline{A}_{R}^{\operatorname{\mathcal{T}}}(x)=\max_{y\in U}\operatorname{\mathcal{T}}(R(x,y),A(y))$ (2)
2.3 Information and decision systems
An information system roughsets is a pair $(U,\operatorname{\mathcal{A}})$, with $U$ a finite, non-empty universe of objects, and $\operatorname{\mathcal{A}}$ a finite, non-empty set of attributes describing these objects. Each attribute $a\in\operatorname{\mathcal{A}}$ has an associated domain of possible values $V_{a}$ and an associated function $a\colon U\to V_{a}$ which maps an object $x\in U$ to its value $a(x)$ for attribute $a$. An attribute is numerical if its value set is a closed real interval. In the remainder of this paper, we will only work in normalised information systems with numerical condition attributes, where all attributes are scaled such that they have unit range, i.e., such that $V_{a}=[0,1]$ for all attributes $a\in\operatorname{\mathcal{A}}$. This scaling step is important because the original range of a condition attribute often has no bearing on the true importance of the feature as a predictor for the class. To this aim, we will preprocess information systems by performing minimum-maximum normalisation, i.e., we replace each original attribute function $a\in\operatorname{\mathcal{A}}$ with the following alternate attribute function $a^{\prime}$:
$a^{\prime}\colon$ $U\to[0,1]$
$x\mapsto\frac{a(x)-\min\{a(u)\mid u\in U\}}{\max\{a(u)\mid u\in U\}-\min\{a(u)\mid u\in U\}}$
In Section 3, we will sometimes look at novel objects that are not yet part of the information system under consideration. In such a case, it is possible that the value for a condition attribute $a$ of that object $v$ is outside of the range observed in the original information system. We solve this by setting $a^{\prime}(v)$ to 1 if $a(v)$ is larger than $\max\{a(u)\mid u\in U\}$, and setting it to 0 if $a(v)$ is smaller than $\min\{a(u)\mid u\in U\}$.
In the following sections, we will assume each the condition attribute $a$ has been normalised, and stop writing $a^{\prime}$.
If there is no inherent order on the values of an attribute, they can be compared by means of a binary fuzzy indiscernibility relation $R_{i}$ defined on $U\times U$. In this paper, we will use the following relation
$R_{i}(x,y)=1-|a(x)-a(y)|$ (3)
However, imagine a case where an attribute represents length, which does have an inherent order. In such cases, we can use a fuzzy dominance relation $R_{d}$, which is a fuzzy $\operatorname{\mathcal{T}}$-preorder relation for a given t-norm $\operatorname{\mathcal{T}}$ that encodes how much the first argument dominates the second. In this paper, we will use the following relation:
$R_{d}(x,y)=\min\left(1-(a(y)-a(x)),1\right)$ (4)
We can compare two objects in $U$ w.r.t. a subset $B$ of $\operatorname{\mathcal{A}}$ using the fuzzy $B$-indiscernibility relation $R_{B}$, defined as
$R_{B}(x,y)=\operatorname{\mathcal{T}}(\underbrace{R_{a}(x,y)}_{a\in B})$ (5)
for each $x$ and $y$ in $U$ and for a given t-norm $\operatorname{\mathcal{T}}$, where $R_{a}$ is the relation used for comparing the values of attribute $a$. In general, $R_{B}$ is a reflexive fuzzy relation.
A decision system $(U,\operatorname{\mathcal{A}}\cup\{d\})$ with a single decision attribute $d$ is an information system which makes a distinction between the condition attributes $\operatorname{\mathcal{A}}$ and the decision attribute $d$, which is not an element of $\operatorname{\mathcal{A}}$. In this paper, the decision attribute is assumed to be categorical, which means that it has a finite, unordered value domain. We will call the values in $V_{d}$, decision classes. For a given object $u$ in $U$, we will identify its decision class $d(u)$ with the set of all objects that have the same value for the decision attribute. Given a subset of attributes $B\subseteq\operatorname{\mathcal{A}}$, the fuzzy $B$-positive region is a fuzzy set in the universe $U$ that contains each object $u$ to the extent that all objects which are approximately similar to $u$ w.r.t. the attributes in $B$, have the same decision value $d$ CORNELIS2010209 :
$\operatorname{\text{POS}}_{B}(u)=\left(\bigcup_{c\in V_{d}}\underline{c}_{R_{B}}^{\operatorname{\mathcal{I}}}\right)(u)$
It is calculated as the membership degree of $u$ to the union of the fuzzy lower approximations of all decision classes of the decision system.
2.4 Rule induction
In the following sections, we will give a short description of some fuzzy and rough rule induction algorithms. RIPPER COHEN1995115 , and its fuzzified version FURIA FURIA , are state-of-the-art (fuzzy) rule induction algorithms that work in two phases. In the first phase, an initial ruleset is developed using a modified version of Incremental Reduced Error Pruning (IREP*) furnkranz1994incremental . Afterwards, in the second phase, an optimization is applied, which calculates possible alternatives for each rule, and selects the best one. To round out the optimization step, IREP* is once again used to fill in any possible gaps left by selecting alternate rules, this optimization step can be repeated multiple times.
Among rough set based rule induction algorithms, MODLEM Stefanowski1998OnRS is generally considered to be the standard option due to its excellent performance. This algorithm was inspired by LEM2 threediscretmodlem . By contrast to the former, it has no need for a discretisation step, selecting the best cut-off values for numerical attributes during the rule generation. MODLEM follows a heuristic strategy for creating an initial rule by choosing sequentially the “best” elementary conditions according to some heuristic criteria. Learning examples that match this rule are removed from consideration. The process is repeated iteratively as long as some learning examples remain uncovered. The resulting set of rules covers all learning examples.
There have been very few attempts at developing rule induction algorithms based on fuzzy rough set theory. A prominent example is QuickRulesJensenCornelis2009 , a greedy fuzzy rough algorithm based on the feature selection algorithm QuickReduct fuzzyroughsets ; jensen2008computational . It greedily selects the most discerning attribute according to its contribution to the positive region. During these calculations, candidate rules are checked and possibly added to a growing ruleset. Recently, the use of the Hamacher t-norm in the context of QuickRules has been explored in hadamacher .
Other work has largely focused on using classical rough set theory to generate fuzzy rulesets Hsiesh2008 ; ShenChouchoulas02 , but mainly ignores the direct use of fuzzy-rough concepts. The induction of gradual decision rules, based on fuzzy-rough hybridization, has been presented in GRECO2006179 . For this approach, new definitions of fuzzy lower and upper approximations were constructed that avoid the use of fuzzy logical connectives altogether. In this approach, decision rules are induced from lower and upper approximations defined for positive and negative relationships between credibility of premises and conclusions. Only the ordinal properties of fuzzy membership degrees are used. Another fuzzy-rough approach to fuzzy rule induction was presented in WangTsangZhaoChen2007 , where fuzzy reducts are employed to generate rules from data. This method also employs a fuzzy-rough feature selection preprocessing step. In JensenShenbook2008 , a fuzzy decision tree algorithm was proposed, based on fuzzy ID3, that incorporates the fuzzy-rough dependency function as a splitting criterion. A fuzzy-rough rule induction method was also proposed in Hong2006 for generating certain and possible rulesets from hierarchical data. This differs from our approach, in which we will only generate a single ruleset, as we will explain in the next section.
3 FRRI: fuzzy-rough rule induction
Let $(U,\operatorname{\mathcal{A}}\cup\{d\})$ be a decision system, where the universe $U$ is a finite set of objects, $\operatorname{\mathcal{A}}=\{a_{1},a_{2},\dots,a_{m}\}$, for a natural number $m$, is the set of (numerical) condition attributes, and $d$ is the (categorical) decision attribute. We will also refer to this decision system as the training set. The condition attributes are normalised (on the training set) to have a range of $[0,1]$, as explained in Section 2.3. The ordering defined on the set of attributes is fixed but arbitrary, as it is often given with real world data. The FRRI algorithm constructs fuzzy decision rules from this decision system, which can be used to summarize the knowledge contained in the data, and to make predictions on the decision attribute of new, unseen objects that are not in the training set. We will define the format of these rules in Section 3.1.
In FRRI, each rule in the final ruleset corresponds to an original object in $U$. Our algorithm consists of two basic steps:
  1. 1.
    Rule shortening: in this step, we look at the initial rules derived from each object of the training set. For each of those rules, we discard conditions, as long as they do not decrease the discerning power of that object. This attribute reduction is performed by constructing a new comparison relation. The shortened rules form an initial ruleset.
  2. 2.
    Rule selection: in this step, we select a minimal number of (shortened) rules from the initial ruleset which still cover the entire training set. This is done by solving an integer programming problem. An optimal solution of this problem defines the final ruleset.
We will now explain each of these steps in detail, but first we will define the rule format used in FRRI.
3.1 Rule format
A rule $r$ in FRRI has the general form
if antecedent then consequent
with the antecedent and the consequent expressed in terms of fuzzy membership. Each rule is originally associated with an object $u_{r}$ from the training set. We will now take a closer look into the exact format of the rule antecedent and consequent.
Definition 3.1.
The antecedent of a rule $r$ is a conjunction of conditions, each of them associated with attribute $a$ in $\operatorname{\mathcal{A}}$. Consider now such attribute $a$ and object $u_{r}$ from $U$. We consider three types of conditions:
  • similar condition: $a$ is similar to the value $a(u_{r})$
  • dominant condition: $a$ is smaller than or similar to the value $a(u_{r})$
  • dominated condition: $a$ is greater than or similar to the value $a(u_{r})$
The antecedent of a rule is then written as a set of (unordered) conditions, linked with and. Note that not all attributes need to be present in a given rule. For the rule $r$, we encode the type of condition associated with each attribute in a type mapping:
[…]
where unused designates that the attribute is not used in that rule.
We look for minimal rules, that is, rules from which no condition attribute can be removed. Moreover, we prefer dominant or dominated conditions over similar, since the former are more general than the latter. The rule shortening step, explained in the next section, is designed to incorporate this preference. The antecedent of a rule induces a fuzzy set, which we can use to calculate the matching degree of an object from the universe to that rule.
Definition 3.2.
Let $r$ be a rule derived from object $u_{r}$. Let $v$ be an object in $U$, and let $\operatorname{\mathcal{T}}$ be a t-norm. The matching degree of $v$ to $r$ is defined as
$M_{r}(v)=\min\left\langle R_{\textsc{type}_{r}(a)}(a(u_{r}),a(v))\mid a\in\operatorname{\mathcal{A}}\right\rangle$
We also denote the fuzzy set $M_{r}$ as the matching set of $r$. The relations corresponding to each condition type are, for $x,y\in[0,1]$:
$R_{\textsc{similar}}(x,y)$ $=R_{i}(x,y)$
$R_{\textsc{dominant}}(x,y)$ $=R_{d}(x,y)$
$R_{\textsc{dominated}}(x,y)$ $=R_{d}(y,x)$
$R_{\textsc{unused}}(x,y)$ $=1$
We can see that dominant and dominated conditions are indeed more general than similar conditions, since, for the same $x\in[0,1]$, it holds that, for any $y\in[0,1]$,
$R_{\textsc{similar}}(x,y)\leq R_{\textsc{dominant}}(x,y)\text{ and }R_{\textsc{similar}}(x,y)\leq R_{\textsc{dominated}}(x,y)$
Definition 3.3.
Let $\operatorname{\mathcal{I}}$ be an implicator. The consequent of a rule $r$ derived from object $u_{r}$, is written as: “$d$ is $d(u_{r})$”. It is encoded as a fuzzy singleton in $V_{d}$:
$\{(d(u_{r}),\underline{d(u_{r})}^{\operatorname{\mathcal{I}}}_{R_{\operatorname{\mathcal{A}}}}(u_{r}))\}$
This membership degree to the lower approximation of its decision class (Equation (1)) represents the degree to which object $u_{r}$ is consistent with its decision class, and can be seen as the confidence score of the rule.
We now have all the information required to define the degree to which a rule covers an object.
Definition 3.4.
Consider a rule $r$ derived from object $u_{r}$ and another object $v$. The rule $r$ covers $v$ to the degree
$S_{r}(v)=\min\left(M_{r}(v),\underline{d(u_{r})}^{\operatorname{\mathcal{I}}}_{R_{\operatorname{\mathcal{A}}}}(u_{r})\right)$
Moreover, we say that the rule $r$ covers object $v$ if $S_{r}(v)>0$, and we call the fuzzy set $S_{r}$ the covering set of $r$.
$a_{1}$ $a_{2}$ $a_{3}$ $a_{4}$ $a_{5}$ $a_{6}$ $a_{7}$ $a_{8}$ $d$
$u_{1}$ 0.12 0.07 0.00 0.04 0.00 0.17 0.45 0.11 0
$u_{2}$ 1.00 0.88 1.00 0.45 0.86 0.52 0.37 1.00 1
$u_{3}$ 0.88 0.60 0.40 0.62 1.00 0.56 0.69 0.56 0
$u_{4}$ 0.88 1.00 0.45 0.55 0.88 0.67 0.10 0.53 1
$u_{5}$ 0.00 0.05 0.95 1.00 0.24 1.00 1.00 0.25 0
$u_{6}$ 0.00 0.11 0.35 0.60 0.35 0.81 0.00 0.00 0
$u_{7}$ 0.12 0.00 0.40 0.00 0.01 0.00 0.20 0.08 0
Table 1: Didactic example used throughout the explanation of this algorithm.
Example 3.5.
Throughout this section, we will use the decision system presented in Table 1 to illustrate the operation of our algorithm. It is a sample of the Pima Indians Diabetes dataset taken from the KEEL dataset repository keeldatarepo , which we have normalised using min-max normalisation, such that all attributes have unit range. We consider a universe with seven objects, eight numerical condition attributes and one binary decision attribute. We consider an example rule $r$ derived from object $u_{1}$ from our example decision system:
if $a_{1}$ is similar to $0.12$ and $a_{2}$ is greater than or similar to $0.07$ and $a_{7}$ is smaller than or similar to $0.45$ then $d$ is 00.
The matching degree of $u_{2}$ to this rule is:
$M_{r}(u_{2})$ $=\min(R_{\textsc{similar}}(0.12,1.00),R_{\textsc{dominated}}(0.07,0.88),R_{\textsc{dominant}}(0.45,0.37))$
$=\min(\begin{aligned} &1-|0.12-1|,\\ &\min\left(1-(0.07-0.88),1\right),\\ &\min\left(1-(0.37-0.45),1\right))\end{aligned}$
$=\min(0.12,1,1)$
$=0.12$
The degree to which this rule covers $u_{2}$ is
$S_{r}(u_{2})=\min(M_{r}(u_{2}),\underline{d(u_{r})}^{\operatorname{\mathcal{I}}}_{R_{\operatorname{\mathcal{A}}}}(u_{r}))=\min(0.12,0.934783)=0.12$
3.2 Rule shortening
As mentioned before, in FRRI, each object of the training set is the starting point of a rule.
Definition 3.6.
Consider an object $u$ in the training set. The total rule $r(u)$ corresponding to $u$ is the rule
if $a_{1}\text{ is similar to }a_{1}(u)\;\textsc{and}$
$a_{2}\text{ is similar to }a_{2}(u)\;\textsc{and}$
$\dots$
$a_{m}\text{ is similar to }a_{m}(u)$
then $d\text{ is }d(u)$
The rule shortening step prunes this total rule of each object to a final shortened, more general rule $\hat{r}(u)$ in the procedure rule_prune. The pseudo-code for this procedure can be found in Algorithm 1. It is repeated independently for each object in the training set. Rule_prune iteratively tries to generalize each condition of $r(u)$ without making the new rule cover objects outside of the class indicated in its consequent. More formally, consider an attribute $a_{i}$ for an object $u$ (line $5$ of Algorithm 1). We consecutively examine the possibility of setting the type of the condition corresponding to $a_{i}$ for the rule corresponding to $u$ as unused, dominant, dominated or similar. If for a type $t$, $S_{r^{*}(u)}$ is a subset of $d(u)$, then we set the type of $a_{i}$ to that type $t$ for this object. If not, we continue to the next type in the list. We repeat this process for each attribute in the order $a_{1},a_{2},\dots,a_{n}$, and we repeat this rule_prune procedure for each object in the training set. The result is a shortened, initial ruleset
^=$\widehat{\mathscr{R}}=\{\hat{r}(u)\mid u\in U\}$
Algorithm 1 The rule_prune procedure
1:  Input: object $u\in U$
2:  Output: shortened rule $\hat{r}(u)$
3:  $t^{*}=(t^{*}_{1},t^{*}_{2},\dots,t^{*}_{m})\leftarrow(\textsc{similar},\textsc{similar},\dots,\textsc{similar})$
4:  for $i=1$ to $m$ do
[…]
6:        $t^{*}_{i}\leftarrow t$
7:        $r^{*}(u)\leftarrow r(u)$ with $t^{*}$ as type vector
8:        if $S_{r^{*}(u)}\subseteq d(u)$ then
9:           break for-loop
10:        end if
11:     end for
12:  end for
13:  return  $\hat{r}(u)\leftarrow r^{*}(u)$
Example 3.7.
Let us apply this rule shortening step to object $u_{1}$ of our example dataset. Consider the first attribute $a_{1}$. We calculate $S_{r^{*}(u_{1})}$ where $r^{*}(u_{1})$ uses the following type mapping:
[…]
We should do this only for those objects from a different class, since for objects $v$ in $d(u_{1})$ it is obvious that
$S_{r^{*}(u_{1})}(v)\leq(d(u_{1}))(v)=1$
We find:
$S_{r^{*}(u_{1})}(u_{2})$ $=\min\left(\min_{2\leq i\leq 8}(R_{\textsc{similar}}(a_{i}(u_{1}),a_{i}(u_{2})),\underline{d(u_{1})}^{\operatorname{\mathcal{I}}}_{R_{\operatorname{\mathcal{A}}}}(u_{1})\right)$
$=0$
$S_{r^{*}(u_{1})}(u_{4})$ $=\min\left(\min_{2\leq i\leq 8}(R_{\textsc{similar}}(a_{i}(u_{1}),a_{i}(u_{4})),\underline{d(u_{1})}^{\operatorname{\mathcal{I}}}_{R_{\operatorname{\mathcal{A}}}}(u_{1})\right)$
$=0$
As such, we can see that $S_{r^{*}(u_{1})}\subseteq d(u_{1})$, and we fix the type of $a_{1}$ for $u_{1}$ as unused. Now we move on to attribute $a_{2}$, and we try the following type mapping:
[…]
Again we calculate $S_{r^{*}(u_{1})}$ with this type mapping, and we find:
$S_{r^{*}(u_{1})}(u_{2})$ $=\min\left(\min_{3\leq i\leq 8}(R_{\textsc{similar}}(a_{i}(u_{1}),a_{i}(u_{2})),\underline{d(u_{1})}^{\operatorname{\mathcal{I}}}_{R_{\operatorname{\mathcal{A}}}}(u_{1})\right)$
$=0$
$S_{r^{*}(u_{1})}(u_{4})$ $=\min\left(\min_{3\leq i\leq 8}(R_{\textsc{similar}}(a_{i}(u_{1}),a_{i}(u_{4})),\underline{d(u_{1})}^{\operatorname{\mathcal{I}}}_{R_{\operatorname{\mathcal{A}}}}(u_{1})\right)$
$=0.058966$
This means we cannot set the type of $a_{2}$ to unused for $u_{1}$. When we calculate $S_{r^{*}(u_{1})}$ with the mapping
[…]
we do find that it is a subset of $d(u_{1})$, and we fix this type as well. We continue this for the remainder of the attributes, until we arrive at the final type mapping of $u_{1}$:
$\textsc{type}_{x_{1}}(a)=\begin{cases}\textsc{dominant}&\text{if }a\in\{a_{2},a_{3}\},\\ \textsc{unused}&\text{otherwise.}\end{cases}$
This results in the following shortened rule $\hat{r}(u_{1})$:
IF $a_{2}$ is smaller than or similar to 0.07 AND $a_{3}$ is smaller than or similar to 00 THEN $d$ is 00.
3.3 Rule selection
After we have finished the rule shortening step, we will select a subset of rules from $\widehat{\mathscr{R}}$ to obtain the final ruleset $\mathscr{R}$. Using Definition 3.4, we introduce the notation $z_{u,v}$ and the binary variable $r_{u}$ for all objects $u,v\in U$:
$z_{u,v}=\begin{cases}1&\text{if }\hat{r}(u)\text{ covers }v\\ 0&\text{otherwise}\end{cases}\;\text{ and }\;r_{u}=\begin{cases}1&\text{if }\hat{r}(u)\in\mathscr{R}\\ 0&\text{otherwise}\end{cases}$
Finally, we can formulate the optimization problem as follows:
[…]
subject to the conditions
$(\forall v\in U)\left(\sum_{u\in U}r_{u}\times z_{u,v}\geq 1\right),$
or in other words: minimize the number of selected rules, while still assuring that every object in the universe is covered by at least one of the selected rules. This integer programming problem can be solved with classical techniques, such as those found in the gurobipy package for Python gurobi .
Algorithm 2 The rule_selection procedure
1:  Input: universe $U$, initial ruleset $\widehat{\mathscr{R}}$
2:  Output: final ruleset $\mathscr{R}$
3:  for all $u$ in $U$ do
4:     for all $v$ in $U$ do
5:        if $S_{\hat{r}(u)}(v)>0$ then
6:           $z_{u,v}\leftarrow 1$
7:        else
8:           $z_{u,v}\leftarrow 0$
9:        end if
10:     end for
11:     define variable $r_{u}$ is 1 iff $\hat{r}(u)\in\mathscr{R}$
12:  end for
13:  Minimize $\sum\limits_{u\in U}r_{u}$ such that $(\forall v\in U)\left(\sum\limits_{u\in U}r_{u}\times z_{u,v}\geq 1\right)$
14:  $\{r_{u}^{*}\mid u\in U\}\leftarrow$ solution of above problem
15:  return  $\{\hat{r}(u)\mid u\in U\text{ and }r_{u}^{*}=1\}$
$u_{1}$ $u_{2}$ $u_{3}$ $u_{4}$ $u_{5}$ $u_{6}$ $u_{7}$
$u_{1}$ 1 0 1 0 0 1 1
$u_{2}$ 0 1 0 1 0 0 0
$u_{3}$ 1 0 1 0 1 0 1
$u_{4}$ 0 1 0 1 0 0 0
$u_{5}$ 1 0 1 0 1 1 1
$u_{6}$ 1 0 1 0 1 1 1
$u_{7}$ 1 0 1 0 1 1 1
Table 2: The variables $z_{u,v}$ for our example dataset.
Example 3.8.
First, we calculate the values of $z_{u,v}$ for our example dataset in Table 2. We can see that no single rule covers the entire training set. However, the following ruleset does:
=$\mathscr{R}=\{\hat{r}(u_{2}),\hat{r}(u_{5})\}$
$\mathscr{R}$ results in a minimal ruleset that still covers the entire training set. The corresponding rules are:
if $a_{3}$ is greater than or similar to 1 and $a_{7}$ is greater than or similar to 0.37 then $d$ is 1
if $a_{1}$ is smaller than or similar to 0 and $a_{2}$ is smaller than or similar to 0.05 then $d$ is 0
In general, the optimal solution is not univocal.
3.4 Inference step
Now, consider a new object $v$ for which we want to predict the decision attribute given its condition attributes and a ruleset $\mathscr{R}$ created by algorithm 2. We derive the inference result as follows. For each possible value $c$ of the decision attribute (i.e., for each class), we calculate the confidence score as the maximal covering degree of $z$ to a rule whose consequent matches class $c$:
$s_{c}=\max\{S_{r}(z)\mid r\in\mathscr{R}\text{ and }d(u_{r})=c\}$
and we assign to $z$ the class with the highest confidence score:
$\arg\max_{c\in V_{d}}s_{c}$
Example 3.9.
Consider a test sample
$v=(0.12,0.40,0.45,0.48,0.00,0.48,0.30,0.50)$
We use the ruleset we found in Example 3.8. The confidence score of class 0 is
$s_{0}$ $=S_{r(u_{5})}(v)$
$=\min(R_{\textsc{dominant}}(0,0.12),R_{\textsc{dominant}}(0.05,0.40),0.945652)$
$=\min(0.78,0.35,0.945652)=0.35$
and the confidence score of class 1 is
$s_{1}$ $=S_{r(u_{5})}(v)$
$=\min(R_{\textsc{dominated}}(1,0.45),R_{\textsc{dominated}}(0.37,0.30),0.600000)$
$=\min(0.55,0.93,0.600000)=0.55$
Therefore, the predicted class of $v$ would be $1$.
4 Experimental evaluation
4.1 Methods and evaluation measures
dataset abbr. $|\operatorname{\mathcal{A}}|$ $|U|$ $|V_{d}|$ IR
australian aus 14 690 2 1.25
bands bands 19 365 2 1.70
bupa bupa 6 345 2 1.38
cleveland cleve 13 297 5 12.62
dermatology derma 34 358 6 5.55
ecoli ecoli 7 336 8 28.60
glass glass 9 214 6 8.44
heart heart 13 270 2 1.25
ionosphere iono 33 351 2 1.79
pima pima 8 768 2 1.87
sonar sonar 60 208 2 1.14
spectfheart spect 44 267 2 38.56
vehicle vehi 18 846 4 1.10
vowel vowel 13 990 11 1.00
wine wine 13 178 3 1.48
winequality-red red 11 1599 6 68.10
wisconsin wisc 9 683 2 1.86
yeast yeast 8 1484 10 92.60
Table 3: Details of the benchmark datasets used for these experiments: abbreviated name, the amount of features of each dataset ($|\operatorname{\mathcal{A}}|$), the number of instances ($|U|$) and classes ($|V_{d}|$), and the imbalance ration IR
We perform a computational experiment to evaluate the performance of our novel FRRI algorithm on 18 benchmark numerical datasets from the KEEL dataset repository [^1] Some of these datasets usually have categorical condition features. We opted to remove these features from the dataset before using them in our experiments. keeldatarepo . The details of these datasets can be found in Table 3, where we list the amount of features of each dataset ($|A|$), as well as the number of instances ($|U|$) and classes ($|V_{d}|$). Finally, the table also contains the imbalance ratio (IR) of each dataset, which is calculated as the size of the largest class divided by the size of the smallest class. We also list an abbreviated name for each dataset.
We implement our algorithm in Python and also compare it with QuickRules JensenCornelis2009 , MODLEM Stefanowski1998OnRS , FURIA FURIA and RIPPER COHEN1995115 . We use our own Python implementation of QuickRules, and we use the implementation found in WEKAWEKA for the other algorithms. We also use the default parameters for all algorithms.
For each algorithm, we perform ten-fold cross-validation on each dataset, using the same folds for each algorithm, after which we calculate the average balanced accuracy of the algorithm on the ten training folds. The balanced accuracy is the arithmetic mean of the recall on each class. Moreover, we also calculate the average number of rules and the average rule length over each ruleset and over each of the ten folds of each dataset. To determine the statistical significance of our results, we apply a two-stage procedure, which starts with a Friedman test multiple-comp-friedman . If the result of that test is significant, we continue with a Conover post-hoc pairwise comparison procedure multiple-comp-conover ; conover_test_big_citations to determine the location of the significant differences. Moreover, we also use the Wilcoxon test pairwise-comp-wilcoxon to detect differences between two methods.
4.2 Results
dataset FRRI MODLEM QuickRules FURIA RIPPER
aus 0.788 0.861 0.865 0.865 0.852
bands 0.675 0.604 0.533 0.629 0.582
bupa 0.601 0.652 0.500 0.657 0.653
cleve 0.290 0.276 0.208 0.245 0.238
derma 0.895 0.941 0.522 0.950 0.942
ecoli 0.721 0.512 0.176 0.530 0.545
glass 0.679 0.518 0.314 0.522 0.495
heart 0.767 0.760 0.786 0.818 0.762
iono 0.912 0.890 0.659 0.887 0.892
pima 0.688 0.699 0.503 0.703 0.699
sonar 0.786 0.705 0.523 0.800 0.705
spect 0.629 0.602 0.559 0.557 0.578
vehicle 0.629 0.740 0.385 0.718 0.688
vowel 0.912 0.684 0.163 0.833 0.705
wine 0.921 0.961 0.916 0.950 0.935
red 0.333 0.339 0.182 0.155 0.150
wisc 0.939 0.955 0.933 0.952 0.953
yeast 0.484 0.454 0.125 0.544 0.550
mean 0.703 0.675 0.492 0.684 0.662
average rank 2.56 2.72 4.44 2.28 3.00
Table 4: Balanced accuracy of FRRI, MODLEM, QuickRules, FURIA and RIPPER on the benchmark datasets.
Table 4 contains the average balanced accuracy of these five algorithms on the benchmark set of datasets, Table 5 contains the average number of rules, and Table 7 the average rule length. In each table, the best result for each dataset is highlighted in bold.
Starting with the balanced accuracy, we can see that FRRI is on average the best performing model, beating the state-of-the-art algorithms on average by a comfortable margin. The Friedman test detected a significant difference ($p<0.001$) between the models. The post-hoc procedure identified that QuickRules performed significantly worse than all other algorithms. We can see that FRRI’s accuracy is never more than $0.11$ lower than the best accuracy on that dataset, while this difference is up to around $0.20$ for MODLEM, FURIA and RIPPER and over $0.74$ for QuickRules. Moreover, of all the algorithms, FRRI is most often the best performer. Finally, when we look at the average rank of each algorithm, we see that FURIA has the lowest average rank, closely followed by FRRI and MODLEM.
Interestingly, when we look at highly imbalanced datasets (IR $>6$), we find that FRRI consistently and significantly ($p<0.05$) outperforms the other algorithms, indicating that our algorithm is able to handle such datasets well.
dataset FRRI MODLEM QuickRules FURIA RIPPER
aus 92 121 732 6 6
bands 59 113 223 19 4
bupa 77 103 332 12 5
cleve 102 95 331 7 4
derma 20 27 90 11 8
ecoli 57 56 315 15 9
glass 51 50 226 13 8
heart 45 62 304 15 4
iono 19 30 460 14 7
pima 125 191 825 13 4
sonar 18 63 276 10 5
spect 24 55 126 17 3
vehicle 162 177 603 31 17
vowel 123 200 1062 69 42
wine 8 13 211 7 5
red 399 376 1541 28 14
wisc 23 31 280 15 5
yeast 485 337 1408 19 21
mean 105 117 519 18 10
Table 5: The average number of rules of FRRI, MODLEM, QuickRules, FURIA and RIPPER on the benchmark datasets.
FRRI MODLEM QuickRules FURIA RIPPER
FRRI 1.00e+0 4.07e-1 9.41e-3 5.13e-2 7.31e-4
MODLEM 4.07e-1 1.00e+0 5.74e-2 6.98e-3 4.03e-5
QuickRules 9.41e-3 5.74e-2 1.00e+0 2.00e-6 3.23e-9
FURIA 5.13e-2 6.98e-3 1.77e-6 1.00e+0 2.44e-1
RIPPER 7.31e-4 4.00e-5 3.23e-9 2.44e-1 1.00e+0
Table 6: The $p$-values obtained by performing the Conover post-hoc pairwise comparison procedure after a significant Friedman test on the ruleset sizes of FRRI, MODLEM, QuickRules, FURIA and RIPPER on the benchmark datasets.
Let us now look at the average size of the rulesets generated by the algorithms, where the Friedman test indicates the presence of significant ($p<0.0001$) differences between the algorithms. Table 6 contains the $p$-values obtained by performing the Conover post-hoc pairwise comparison procedure. First, we notice that FURIA and RIPPER generate significantly smaller rulesets than the (fuzzy)-rough rule induction algorithms. This is because both of these models use an iterative pruning process, where old rules are repeatedly replaced with more general versions. Moreover, QuickRules is by far the worst performer. One reason this might be the case, is that the iterative process of QuickRules considers the objects in the training set many times, and looks for a total ruleset that covers the entire dataset in a greedy way. Without any pruning, this is sure to deliver much more rules that algorithms with better heuristics (MODLEM), pruning (FURIA and RIPPER) and those that use the exact solutions of an optimization problem (FRRI). Finally, let us zoom in on the contrast between FRRI and MODLEM, which is even more pronounced here. FRRI generates the smaller ruleset of the two in all but 5 of our datasets. It is interesting to observe that those datasets are highly imbalanced. Actually, this maybe considered further evidence that FRRI is better suited towards dealing with imbalanced data, as it will generate enough rules to cover both the smallest and biggest classes completely. For MODLEM, the worst case is vowel, with where it needs 77 more rules. On the other hand, statistical testing again shows that both MODLEM and FRRI are significantly better ($p<0.001$) than QuickRules. Using a one-sided Wilcoxon test, we can establish that FRRI is also significantly better ($p<0.05$) than MODLEM w.r.t. the size of the induced ruleset.
dataset FRRI MODLEM QuickRules FURIA RIPPER
aus 5.15 $\pm$ 1.5 2.36 $\pm$ 0.9 8.05 $\pm$ 1.9 2.17 $\pm$ 0.7 2.33 $\pm$ 0.7
bands 6.29 $\pm$ 2.0 2.11 $\pm$ 0.6 1.75 $\pm$ 0.6 5.58 $\pm$ 1.4 4.75 $\pm$ 1.1
bupa 4.20 $\pm$ 1.0 2.19 $\pm$ 0.6 4.40 $\pm$ 1.0 3.17 $\pm$ 0.8 2.40 $\pm$ 0.8
cleve 5.45 $\pm$ 1.4 2.59 $\pm$ 1.0 8.08 $\pm$ 2.0 3.86 $\pm$ 0.8 2.50 $\pm$ 0.9
derma 4.70 $\pm$ 2.1 2.56 $\pm$ 1.0 6.73 $\pm$ 3.5 2.73 $\pm$ 1.1 1.75 $\pm$ 0.8
ecoli 3.85 $\pm$ 0.9 2.34 $\pm$ 0.8 3.82 $\pm$ 1.0 3.00 $\pm$ 0.8 2.00 $\pm$ 0.8
glass 3.88 $\pm$ 1.1 2.10 $\pm$ 0.8 5.34 $\pm$ 1.5 3.38 $\pm$ 1.1 2.38 $\pm$ 0.9
heart 4.90 $\pm$ 1.3 2.21 $\pm$ 0.7 7.58 $\pm$ 1.9 3.33 $\pm$ 1.0 1.50 $\pm$ 0.5
iono 4.53 $\pm$ 2.2 1.53 $\pm$ 0.6 11.7 $\pm$ 5.4 2.71 $\pm$ 1.4 1.43 $\pm$ 0.5
pima 5.08 $\pm$ 1.2 2.07 $\pm$ 0.7 5.69 $\pm$ 1.3 2.69 $\pm$ 1.0 2.25 $\pm$ 0.8
sonar 7.80 $\pm$ 2.0 1.56 $\pm$ 0.5 21.5 $\pm$ 10 6.20 $\pm$ 1.4 4.60 $\pm$ 2.3
spect 8.24 $\pm$ 3.2 1.75 $\pm$ 0.5 1.17 $\pm$ 0.4 3.06 $\pm$ 1.1 3.00 $\pm$ 1.4
vehi 5.51 $\pm$ 1.4 2.60 $\pm$ 0.7 1.54 $\pm$ 0.5 3.81 $\pm$ 1.1 2.59 $\pm$ 0.8
vowel 6.22 $\pm$ 1.5 2.15 $\pm$ 0.9 6.49 $\pm$ 1.9 3.74 $\pm$ 1.4 3.19 $\pm$ 1.3
wine 4.51 $\pm$ 1.4 1.46 $\pm$ 0.5 6.97 $\pm$ 2.4 2.14 $\pm$ 0.6 1.60 $\pm$ 0.5
red 5.84 $\pm$ 1.4 2.81 $\pm$ 0.9 6.76 $\pm$ 1.7 4.68 $\pm$ 1.3 3.43 $\pm$ 1.3
wisc 3.74 $\pm$ 1.2 2.23 $\pm$ 0.7 3.84 $\pm$ 1.1 3.47 $\pm$ 0.7 1.80 $\pm$ 0.4
yeast 4.82 $\pm$ 1.0 3.26 $\pm$ 1.2 4.93 $\pm$ 0.9 3.11 $\pm$ 0.9 2.81 $\pm$ 1.1
mean 5.26 2.21 6.46 3.49 2.57
Table 7: The mean and the standard deviation of rule lengths in the rulesets of FRRI, MODLEM, QuickRules, FURIA and RIPPER on the benchmark datasets.
Finally, we consider the lengths of the generated rules. Here, FRRI still performs better than QuickRules, but not significantly so, unlike MODLEM, FURIA and RIPPER, which induce significantly ($p<0.01$) shorter rules than both QuickRules and FRRI. We examine the distribution of rule lengths across the benchmark datasets for FRRI. We find that it does not generate a few, long hyperspecific rules, which would have allowed us to easily decrease the average length by removing those rules. However, FRRI does produce rulesets with more varied rule lengths than the other algorithms, except for QuickRules, which does tend to generate long rules.
5 Conclusions and future work
In this article, we introduced a novel rule induction algorithm, Fuzzy-Rough Rule Induction (FRRI), which combines the best parts of fuzzy and rough set theory to induce small rulesets consisting of high-accuracy rules. Our algorithm starts by discarding unnecessary information from the objects in our dataset, and then selects a minimal amount of objects to serve as rules. These rules can be viewed as summaries of the data, which are put to work to classify new samples. Experimental evaluation of the performance of FRRI against state-of-the-art rule induction algorithms in this last task showed that our algorithm is more accurate while using fewer rules. Moreover, the obtained rules are not much longer than those induced by state-of-the-art rough set algorithms.
We have identified some challenges for future work:
  1. 1.
    We want to adapt our algorithm to regression and ordinal classification problems.
  2. 2.
    We will explore the impact of ordering the attributes during the rule shortening phase. This can be either a single pass before the phase itself, or applying a different order for each object. Various strategies could be used, like randomizing the order of attributes, or using a discriminative ordering. Moreover, we could also use all possible minimal rules that could be created from a single object.
  3. 3.
    In this version of FRRI, we were aiming at obtaining the minimal amount of rules, which is why we solved the optimisation problem exactly in Section 3.3. Since this formulation of the problem is equivalent to the vertex cover problem from graph theory, we know it is NP-hard and cannot be solved optimally with lower complexity. However, to reduce time complexity when applying FRRI to big data, we want to explore approximate solvers, taking inspiration from the work done on approximate solvers for the vertex cover problem. Additionally, when looking at big data, we are interested in exploring a bagging ensemble approach, where we generate smaller rulesets on subsets of the data and creating one ensemble ruleset.
  4. 4.
    We also want to construct hierarchical rulesets, by combining similar rules into a higher-level, more general rule, thus reducing the amount of rules and improving the explainablility.
Acknowledgements
Henri Bollaert would like to thank the Special Research Fund of Ghent University (BOF-UGent) for funding his research.